# A unified concept of the degree of ill-posedness for compact and non-compact linear operators in Hilbert spaces under the auspices of the spectral theorem

### v1

@@ 3.3 Ill-posedness of multiplication operators on the real axis
    (a1) $\;\lambda(\omega)=\frac{1}{(1+\omega^{2})^{s}}\quad(\omega\in\mathbb{R},\;s>0)\quad\mbox{implies}\quad\Phi_{\lambda,\mu_{\mathrm{Leb}}}(\epsilon)=2\mu(\{\omega\geq 0:\,1+\omega^{2}\leq\epsilon^{-1/s}\})\,,$ hence $\;\Phi_{\lambda,\mu}(\epsilon)=2\sqrt{\epsilon^{-\frac{1}{s}}-1}\asymp\epsilon^{-\frac{1}{2s}}$ as $\epsilon\to 0$. The corresponding interval (9) hence reduces to a singleton $\left[A_{\lambda,\mu_{\mathrm{Leb}}},B_{\lambda,\mu_{\mathrm{Leb}}}\right]=\{s\}$, indicating a moderate degree $s>0$ of ill-posedness for the underlying operator equation here.
@@
    The impact of this inner zero on the asymptotics of $\Phi_{\lambda,\mu_{\mathrm{Leb}}}(\epsilon)$ as $\epsilon\to 0$ is negligible (cf. also [26, Prop. 1]), and due to the zeros of $\lambda(\omega)$ at infinity $\omega\to\pm\infty$ we have $\Phi_{\lambda,\mu_{\mathrm{Leb}}}(\epsilon)\asymp\epsilon^{-\frac{1}{2}}$ as well as $\left[A_{\lambda,\mu_{\mathrm{Leb}}},B_{\lambda,\mu_{\mathrm{Leb}}}\right]=\{1\}$, expressing a degree one of ill-posedness.
@@
3.4 Ill-posedness concept covering compact and non-compact operators
Covering ill-posed problems with compact and non-compact operators regarding the nature and degree of ill-posedness is a never ending story written by many authors in the inverse problems literature. For example, we refer to the claim of Nashed in [27] that non-degenerating compact operators mapping between infinite dimensional Hilbert spaces are always more ill-posed than the non-compact ones. Arguments for that claim come from the comparison of ill-posed situations with two linear operators by partial orderings and associated range inclusions (see, e.g, [16]). In this context, the range of non-compact operator can never be a subset of the range of a compact operator. Also in [16], however, it was shown by observing the modulus of injectivity for various discretizations that the ill-posedness nature of non-compact operators is completely different from the nature of compact ones and that hence the simple claim of Nashed is rather problematic, because sequences of compact operators do never converge in norm to a non-compact operator. In the present paper, we try to add some new facets to the story of covering compact and non-compact cases with respect to the degree of ill-posedness under the auspices of the spectral theorem, where we do not compare two operators, but whole families of operators possessing comparable properties. Following this approach, the same degree of ill-posedness may occur for classes of compact operators and classes of non-compact operators.
in $\Omega$, because the functions $\xi,\zeta\in L^{2}(\Omega,\mu)$ appear by unitary transformation from $u$ and $h$.
In order to overcome the difficulties arising from the ambiguity of possible measure spaces $(\Omega,\mu)$, and inspired by the examples from the previous subsection, we will choose for the case $\mu(\Omega)=\infty$ a benchmark measure space with focus on the Lebesgue measure $\mu_{\mathrm{Leb}}$:
@@ 3.4 Ill-posedness concept covering compact and non-compact operators
$T:\mathcal{X}\to Y$ is a bounded linear operator between infinite-dimensional Hilbert spaces $\mathcal{X}$ and $\mathcal{Y}$ such that there exists a spectral decomposition (5) with $\Omega\subseteq\mathbb{R}^{d}$ and $\mu=\mu_{\mathrm{Leb}}$.
@@
This assumption might be restrictive at the moment, but we will discuss several examples satisfying this assumption below, e.g. convolution operators and the Hausdorff moment problem. We furthermore emphasize that this assumption is natural in the sense that also for the compact case a benchmark measure space (namely $\Omega=\mathbb{N}$ and the counting measure $\mu_{\mathrm{disc}}$) has been chosen. Taking note of the transformations as in Lemma 1, the restriction to a situation with Lebesgue measure has the rationale background that its translation invariance ensures all frequencies to have the same influence. In the following we write
$\Phi_{\lambda}:=\Phi_{\lambda,\mu_{\mathrm{Leb}}}$
under Assumption 2.
Definition 2.
Let $\mathcal{X}$ and $\mathcal{Y}$ denote infinite-dimensional Hilbert spaces. Moreover, let $T:\mathcal{X}\to\mathcal{Y}$ be either an injective and compact linear operator, or alternatively an injective and bounded non-compact linear operator with non-closed range, each mapping between $\mathcal{X}$ and $\mathcal{Y}$ such that the Assumptions 1 and 2 are satisfied w.r.t. the distribution function $\Phi_{\lambda}$ from (7) and for the benchmark Lebesgue measure $\mu_{\mathrm{Leb}}$. Then the interval of ill-posedness of $T$ is defined as
$\left[A_{\lambda},B_{\lambda}\right]:=\left[\liminf\limits_{\epsilon\searrow 0}\frac{\log(\epsilon)}{-2\log\left(\Phi_{\lambda}(\epsilon)\right)},\limsup\limits_{\epsilon\searrow 0}\frac{\log(\epsilon)}{-2\log\left(\Phi_{\lambda}(\epsilon)\right)}\right],$
and the operator equation (1) is called mildly ill-posed] whenever $A_{\lambda}=B_{\lambda}=0$, severely ill-posed whenever $A_{\lambda}=B_{\lambda}=\infty$, moderately ill-posed whenever $0<A_{\lambda}\leq B_{\lambda}<\infty$, and ill-posed of degree $s>0$], if $A_{\lambda}=B_{\lambda}=s$.
Theorem 2.
For compact operators $T$, the degrees of ill-posedness from Definitions 1 and 2 agree.
@@
If $T$ is compact, we can rewrite the counting function (3) as distribution function (7) with $\Omega=[0,\infty)$ and Lebesgue measure $\mu=\mu_{\mathrm{Leb}}$ on $\mathbb{R}$ in the form
$\Phi_{\lambda}(\epsilon)=\mu_{\mathrm{Leb}}(\{\omega\in[0,\infty):\,\lambda(\omega)>\epsilon)\})\quad(\epsilon>0),$
where the piece-wise constant multiplier function $\lambda$ attains the form
$\lambda(\omega)=\{\sigma_{n}^{2}\;\mbox{if}\;n-1\leq\omega<n\;(n=1,2,...)\}\quad(\omega\in[0,\infty).$
This way, we have found a spectral representation of $T^{*}T$ where $\mu=\mu_{\mathrm{Leb}}$ is the Lebesgue-measure, $\Omega=[0,\infty)$, and hence Assumptions 1(a) and 2 are satisfied. Furthermore, since the counting function (3) agrees with the distribution function $\Phi_{\lambda}$, Assumption 1(b) is fulfilled as well and both definitions of the degree of ill-posedness agree. ∎
Let us briefly comment on the decreasing rearrangement of the multiplier function:
@@
Under Assumptions 1 and 2 and in particular due to $\mu_{\mathrm{Leb}}(\Omega)=\infty$, it makes sense to characterize the degree of ill-posedness by considering the behavior of the multiplier function $\lambda$ at infinity. As already suggested in [26], the decreasing rearrangement $\lambda^{*}$ of $\lambda$ defined by the (generalized) inverse function
$\lambda^{*}(t):=\inf\left\{\tau>0:\,\Phi_{\lambda}(\tau)\leq t\right\}\quad(0\leq t<\infty)$ (11)
of $\Phi_{\lambda}$ seems to be an appropriate tool for that. Precisely, we have that ${\rm essran}(\lambda)={\rm essran}(\lambda^{*})={\rm spec}(T^{*}T)$. Furthermore, $\lambda^{*}(0)=\|T\|^{2}$ as a consequence of $\Phi_{\lambda}(\epsilon)=0$ for $\epsilon\geq\|T\|^{2}$ and $\lambda^{*}(t)$ tends monotonically to zero as $t\to\infty$. The well-defined decay rate of $\lambda^{*}$ to zero at infinity can be used for measuring the degree of ill-posedness, because $\lambda$ and $\lambda^{*}$ are spectrally equivalent, which means that $\Phi_{\lambda}$ is distribution function for both $\lambda$ and $\lambda^{*}$.
@@ 4.1 The Hausdorff moment operator
$\Phi_{\lambda}(\epsilon)\approx\frac{1}{\pi}\log\left(\frac{2\pi}{\epsilon}\right)\quad\mbox{for sufficiently small}\;\epsilon>0.$
This gives as in case (b) of Example 7 that $[A_{\lambda},B_{\lambda}]=\{\infty\}$, which indicates severe ill-posedness of the Hausdorff moment problem.
@@ 4.2 Convolution operators
with a convolution kernel $h\in L^{1}\left(\mathbb{R}^{d}\right)$. Injectivity of $T$ requires with the Lebesgue measure $\mu_{\mathrm{Leb}}$ in $\mathbb{R}^{d}$ the condition
@@
$\mathcal{F}\,T^{*}T\mathcal{F}^{-1}=M_{|\mathcal{F}h|^{2}}\,,$
where the associated measure space is formed by $\Omega=\mathbb{R}^{d}$ with the Lebesgue measure $\mu_{\mathrm{Leb}}$ such that the case $\mu_{\mathrm{Leb}}(\Omega)=\infty$ applies. For $h\in L^{1}\left(\mathbb{R}^{d}\right)$, we have that the complex function $[\mathcal{F}h](\xi)\;(\xi\in\mathbb{R}^{d})$ is bounded, continuous and tends to zero as $\|\xi\|$ tends to infinity. Thus, also the non-negative multiplier function $\lambda(\omega)=|[\mathcal{F}h](\omega)|^{2}\;(\omega\in\mathbb{R}^{d})$ is bounded, continuous and tends to zero as $\|\omega\|\to\infty$. This indicates for $h\not=0$, which is excluded by the injectivity of $T$, pure continuous spectrum of $T^{*}T$, non-compactness of this operator and, as a consequence of $0\in{\rm essran}(\lambda)={\rm spec}(T^{*}T)$, ill-posedness of type I in the sense of Nashed of the underlying linear integral equation (1) of the first kind of convolution type in the Hilbert space $L^{2}\left(\mathbb{R}^{d}\right)$.
@@
The linear convolution operator (15) mapping in $L^{2}\left(\mathbb{R}^{d}\right)$ is well-defined and bounded as soon as $h\in L^{1}\left(\mathbb{R}^{d}\right)$. Moreover, the multiplier function $\lambda$ assigned to $T^{*}T$ for a measure space with $\Omega=\mathbb{R}^{d}$ and the Lebesgue measure $\mu_{\mathrm{Leb}}$ in $\mathbb{R}^{d}$ satisfies Assumption 1 whenever $h\in L^{p}\left(\mathbb{R}^{d}\right)$ for some $1<p\leq 2$.
@@
(see, e.g, [2, Theorem 1]). This implies for $h\in L^{p}\left(\mathbb{R}^{d}\right)\;(1<p\leq 2)$ that we have $\lambda=|\mathcal{F}h|^{2}\in L^{r}(\mathbb{R}^{d})$ with $r=\frac{p}{2(p-1)}\in[1,\infty)$. Consequently, Corollary 2 is applicable. This completes the proof. ∎
@@
$\Phi_{\lambda}(\epsilon)=\mu_{\mathrm{Leb}}\left(\omega\in\mathbb{R}^{d}:\,\|\omega\|<\sqrt{2\log(\pi^{d}/\epsilon)}\right)=\frac{\pi^{d/2}}{\Gamma(d/2)\,d}\,\left(\sqrt{2\log(\pi^{d}/\epsilon)}\right)^{d}$
@@
$\Phi_{\lambda}(\epsilon)\asymp\;\left[\log(1/\epsilon)\right]^{\frac{d}{2}}\quad\mbox{as}\quad\epsilon\searrow 0,$
which shows $A_{\lambda}=B_{\lambda}=\infty$ indicating severe ill-posedness.
@@
$\Phi_{\lambda}(\epsilon)\asymp\epsilon^{-\frac{1}{4a}}\quad\mbox{as}\quad\epsilon\searrow 0,$

### v2

@@ 3.3 Ill-posedness of equations with multiplication operators on the real axis
    (a1) $\;\lambda(\omega)=\frac{1}{(1+\omega^{2})^{s}}\quad(\omega\in\mathbb{R},\;s>0)\quad\mbox{implies}\quad\Phi_{\lambda,\mu_{\mathrm{Leb}}}(\epsilon)=2\mu(\{\omega\geq 0:\,1+\omega^{2}\leq\epsilon^{-1/s}\})\,,$ hence $\;\Phi_{\lambda,\mu_{Leb}}(\epsilon)=2\sqrt{\epsilon^{-\frac{1}{s}}-1}\asymp\epsilon^{-\frac{1}{2s}}$ as $\epsilon\to 0$. The corresponding interval (8) hence reduces to a singleton $\left[A_{\lambda,\mu_{\mathrm{Leb}}},B_{\lambda,\mu_{\mathrm{Leb}}}\right]=\{s\}$, indicating a moderate degree $s>0$ of ill-posedness for the underlying operator equation here.
@@
    The impact of this inner zero on the asymptotics of $\Phi_{\lambda,\mu_{\mathrm{Leb}}}(\epsilon)$ as $\epsilon\to 0$ is negligible (cf. also [28, Prop. 1]), and due to the zeros of $\lambda(\omega)$ at infinity $\omega\to\pm\infty$ we have $\Phi_{\lambda,\mu_{\mathrm{Leb}}}(\epsilon)\asymp\epsilon^{-\frac{1}{2}}$ as well as $\left[A_{\lambda,\mu_{\mathrm{Leb}}},B_{\lambda,\mu_{\mathrm{Leb}}}\right]=\{1\}$, expressing a degree one of ill-posedness.
@@
3.4 Chances and limitations of Lebesgue measure spaces
Lemma 1 has shown that the strength of ill-posedness expressed by interval and degree of ill-posedness in the sense of Definitions 1 and 2 strongly depends on the choice of the measure space $(\Omega,\mu)$. In order to overcome the difficulties arising from the ambiguity of possible measure spaces, we have great expectations of benchmark measure spaces. By using such benchmarks, the outcome of Definition 2 is uniquely determined for the operator $T$. The use of $(\mathbb{N},\mu_{disc})$ with the counting measure $\mu_{disc}$ for the classification of compact operators is fairly uncontroversial. Likewise, the examples in Subsection 3.3 establish the Lebesgue measure $\mu_{Leb}$ on $\mathbb{R}$ as a benchmark measure for the classification of families of non-compact operators, which motivates the following Assumption 2 as a supplement to Assumption 1.
@@ 3.4 Chances and limitations of Lebesgue measure spaces
Let $T:\mathcal{X}\to Y$ be a injective bounded non-compact linear operator with non-closed range mapping between infinite-dimensional Hilbert spaces $\mathcal{X}$ and $\mathcal{Y}$ such that there exists a spectral decomposition (5) with a convex set $\Omega\subseteq\mathbb{R}^{d}$, $\mu(\Omega)=\infty$ and $\mu=\mu_{\mathrm{Leb}}$.
@@
This assumption might be restrictive at the moment, but we will discuss below in Section 4 several examples with non-compact operators satisfying this assumption, e.g. convolution operators and the Hausdorff moment problem. Let us emphasize that the Lebesgue measure employed in Assumption 2 is natural. Taking note of the transformations as in Lemma 1, the restriction to a situation with Lebesgue measure has the rationale background that its translation invariance ensures all frequencies to have the same influence.
It would be desirable to find a common measure space $(\Omega,\mu)$ which allows for a comparison of the ill-posedness strength between compact and non-compact operator. For example, an open question seems to be whether a severely ill-posed non-compact operator exists which in any sense proves to be more ill-posed than another moderately ill-posed compact operator. Trying to extend Assumption 2 to compact cases, it had been an idea to use $\Omega=[0,\infty)$ with the Lebesgue measure $\mu_{Leb}$ over $\mathbb{R}$ as possible common measure space by exploiting the (generalized) inverse
$\lambda^{*}(\omega):=\inf\left\{\tau>0:\,\Phi_{\lambda,\mu}(\tau)\leq\omega\right\}\quad(0\leq\omega<\infty)$ (10)
of the distribution function $\Phi_{\lambda,\mu}$. This decreasing rearrangement of the multiplier function $\lambda$ is defined on this subset $[0,\infty)$ of the real axis, even if $\lambda$ is originally defined on a subset of $\mathbb{R}^{d}$ for $d>1$. Moreover, $\lambda$ and $\lambda^{*}$ are closely related, because they have the same distribution function $\Phi_{\lambda,\mu}(\epsilon)$, and the growth rate of this distribution functions as $\epsilon\searrow 0$, which is relevant for Definition 2, can be rewritten under Assumption 1 as decay rate of $\lambda^{*}(t)$ as $t\to\infty$, see in this context also [28].
Following this idea, compact operators $T=T_{comp}$ with singular values $\{\sigma_{n}\}_{n=1}^{\infty}$ and distribution (counting) function $\Phi$ from (3) possess the non-negative, non-increasing and piece-wise constant multiplier function $\lambda_{comp}$ defined by formula
$\lambda_{comp}(\omega)=\{\sigma_{n}^{2}\;\mbox{if}\;n-1\leq\omega<n\;(n=1,2,...)\}\quad(0\leq\omega<\infty)$ (11)
as inverse of $\Phi(\epsilon)$ tending to zero as $\omega\to\infty$. This function from (11) is the decreasing rearrangement of the original multiplier function $\lambda_{comp}(n)=\sigma_{n}^{2}\;(n\in\mathbb{N})$ assigned to $T_{comp}^{*}T_{comp}$. By bringing the Lebesgue measure over $\mathbb{R}$ into play, we find that
$\Phi_{\lambda_{comp},\mu_{Leb}}(\epsilon)=\mu_{\mathrm{Leb}}(\{\omega\in[0,\infty):\,\lambda_{comp}(\omega)>\epsilon\})\quad(\epsilon>0)$
coincides with the corresponding counting function (3). If we consider the corresponding multiplication operator $M_{\lambda}$ mapping in $L^{2}([0,\infty),\mu_{Leb})$ defined by formula (6) for the multiplier function $\lambda=\lambda_{comp}$ from (11), then this non-negative self-adjoint multiplication operator $M_{\lambda_{comp}}$ has a pure point spectrum, and the eigenvalues of $M_{\lambda_{comp}}$ are the same values as those of the compact operator $T_{comp}^{*}T_{comp}$. Namely, the essential range of the multiplier function and the spectrum values for a multiplication operator are identical. However, the following general assertion of Proposition 2 shows the limitation of the Lebesgue measure in this setting. Here, $M_{\lambda_{comp}}$ is a non-compact operator with infinite dimensional eigenspaces for all eigenvalues, and the spectrum of $M_{\lambda_{comp}}$ is not a discrete spectrum with only finite dimensional eigenspaces, which would be necessary for being a compact self-adjoint operator like $T_{comp}^{*}T_{comp}$. As a consequence we have that $M_{\lambda_{comp}}$ as a non-compact operator cannot be derived from the compact operator $T_{comp}^{*}T_{comp}$ by a unitary transformation, This implies that a desired comparison of compact and non-compact operators fails on this way.
Proposition 2.
There is no compact injective non-negative multiplication operator
$M_{\lambda}:L^{2}([0,\infty),\mu_{Leb})\to L^{2}([0,\infty),\mu_{Leb})$
with real-valued multiplier function $\lambda\in L^{\infty}([0,\infty),\mu_{Leb})$.
@@
First note, that $M_{\lambda}$ is self-adjoint for any real-valued $\lambda$. Furthermore, it is bounded for $\lambda\in L^{\infty}([0,\infty),\mu_{Leb})$.
Since $M_{\lambda}$ is injective and non-negative, the function $\lambda$ must be non-zero and non-negative itself. Hence, there exists a $c>0$ such that the set $S:=\{\lambda\geq c\}$ satisfies $\mu_{Leb}\left(S\right)>0$. On the Hilbert space of all $L^{2}$-functions supported in $S$, i.e.
$\mathcal{V}:=\left\{f\in L^{2}([0,\infty),\mu_{Leb})\mid f=f\cdot\chi_{S}\right\},$
the operator $M_{\lambda}$ is surjective (since for $h\in\mathcal{V}$ it holds $M_{\lambda}\left(h\chi_{S}\lambda^{-1}\right)=h\chi_{S}=h$) and bounded (since $\chi_{S}\lambda^{-1}\leq c^{-1}<\infty$). So if $M_{\lambda}:L^{2}([0,\infty),\mu_{Leb})\to L^{2}([0,\infty),\mu_{Leb})$ was compact, so was
$\text{id}_{\mathcal{V}}=\left(\left(M_{\lambda}\right)_{|_{\mathcal{V}}}\right)^{{}_{1}}\left(M_{\lambda}\right)_{|_{\mathcal{V}}}:\mathcal{V}\to\mathcal{V}.$
However, this is impossible, since $\mathcal{V}$ is infinite dimensional: There exists a partition $S=\bigcup_{i\in\mathbb{N}}S_{i}$ with pairwise disjoint sets $S_{i}$, $\mu_{Leb}(S_{i})>0$ for all $i\in\mathbb{N}$, the indicator functions of which are linearly independent. ∎
@@
Note, that we exploited a specific property of the Lebesgue-measure here, namely that for each set $S\subseteq\mathbb{R}$ with $\mu_{Leb}(S)>0$ there exists an infinite partition $S=\bigcup_{i\in\mathbb{N}}S_{i}$ of pairwise disjoint sets $S_{i}$ with $\mu_{Leb}(S_{i})>0$ for all $i\in\mathbb{N}$. However, this property is also true for any measure without atoms, including e.g. the Cantor measure on $[0,1]$.
@@ 4.1 The Hausdorff moment operator
$\Phi_{\lambda,\mu_{Leb}}(\epsilon)\approx\frac{1}{\pi}\log\left(\frac{2\pi}{\epsilon}\right)\quad\mbox{for sufficiently small}\;\epsilon>0.$ (15)
Evidently, Assumptions 1 and 2 agree in that example, and formula (15) gives as in case (b) of Example 7 that $[A_{\Phi_{\lambda,\mu_{Leb}}},B_{\Phi_{\lambda,\mu_{Leb}}}]=\{\infty\}$, which indicates severe ill-posedness of the Hausdorff moment problem.
@@ 4.2 Convolution operators
with a convolution kernel $h\in L^{1}\left(\mathbb{R}^{d}\right)$. The measure space $(\Omega,\mu)$ is here $(\mathbb{R}^{d},\mu_{Leb})$ with the multi-dimensional Lebesgue measure over $\mathbb{R}^{d}\;(d\in\mathbb{N})$ and $\mu_{Leb}(\Omega)=\infty$. Injectivity of $T$ requires the condition
@@
$\mathcal{F}\,T^{*}T\mathcal{F}^{-1}=M_{|\mathcal{F}h|^{2}}\,.$
For $h\in L^{1}\left(\mathbb{R}^{d}\right)$, we have that the complex function $[\mathcal{F}h](\xi)\;(\xi\in\mathbb{R}^{d})$ is bounded, continuous and tends to zero as $\|\xi\|$ tends to infinity. Thus, also the non-negative multiplier function $\lambda(\omega)=|[\mathcal{F}h](\omega)|^{2}\;(\omega\in\mathbb{R}^{d})$ is bounded, continuous and tends to zero as $\|\omega\|\to\infty$. This indicates for $h\not=0$, which is excluded by the injectivity of $T$, pure continuous spectrum of $T^{*}T$, non-compactness of this operator and, as a consequence of $0\in{\rm essran}(\lambda)={\rm spec}(T^{*}T)$, ill-posedness of type I in the sense of Nashed of the underlying linear integral equation (1) of the first kind of convolution type in the Hilbert space $L^{2}\left(\mathbb{R}^{d}\right)$.
@@
The linear convolution operator (16) mapping in $L^{2}\left(\mathbb{R}^{d}\right)$ is well-defined and bounded as soon as $h\in L^{1}\left(\mathbb{R}^{d}\right)$. Moreover, the multiplier function $\lambda$ assigned to $T^{*}T$ for a measure space with $\Omega=\mathbb{R}^{d}$ and the Lebesgue measure $\mu_{\mathrm{Leb}}$ over $\mathbb{R}^{d}$ satisfies Assumptions 1 and 2 whenever $h\in L^{p}\left(\mathbb{R}^{d}\right)$ for some $1<p\leq 2$.
@@
(see, e.g, [2, Theorem 1]). This implies for $h\in L^{p}\left(\mathbb{R}^{d}\right)\;(1<p\leq 2)$ that we have $\lambda=|\mathcal{F}h|^{2}\in L^{r}(\mathbb{R}^{d})$ with $r=\frac{p}{2(p-1)}\in[1,\infty)$. Consequently, Corollary 2 is applicable. The proof is complete, because Assumption 2 is trivially satisfied. ∎
@@
$\Phi_{\lambda,\mu_{Leb}}(\epsilon)=\mu_{\mathrm{Leb}}\left(\omega\in\mathbb{R}^{d}:\,\|\omega\|<\sqrt{2\log(\pi^{d}/\epsilon)}\right)=\frac{\pi^{d/2}}{\Gamma(d/2)\,d}\,\left(\sqrt{2\log(\pi^{d}/\epsilon)}\right)^{d}$
@@
$\Phi_{\lambda,\mu_{Leb}}(\epsilon)\asymp\;\left[\log(1/\epsilon)\right]^{\frac{d}{2}}\quad\mbox{as}\quad\epsilon\searrow 0,$
which shows $A_{\Phi_{\lambda,\mu_{Leb}}}=B_{\Phi_{\lambda,\mu_{Leb}}}=\infty$ indicating severe ill-posedness.
@@
$\Phi_{\lambda,\mu_{Leb}}(\epsilon)\asymp\epsilon^{-\frac{1}{4a}}\quad\mbox{as}\quad\epsilon\searrow 0,$